1. An Employee at Laser Kinetics borrows $2,000 on September 1 and must repay a total of S 2,800 exactly 1 year later. Determine the interest amount and the interest rate paid. 2. a) Calculate the amount deposited 1 year ago to have $ 750 at an interest rate of 25% per year b) Calculate the amount of interest eamed during this time period.
Answer #1

● 1. Money borrowed by an employee on September 1 = $2,000

Money to be repaid exactly 1 year later = $2,800.

(i) Interest amount = $800 ($2,800-2,000).

(ii) Interest rate calculation:

A = P(1+i)^n

2,800 = 2,000(1+i)^1

1+i = 2,800/2,000 = 1.4

i. = 1.4 – 1 = 0.4

Interest rate = 40%.

2. a) Amount earned after 1 year = $750.

Interest rate = 25%

Amount deposited one year ago:

P = A/(1+i)^1

P = 750/1.25

Amount deposited one year ago (P) = $600.

ii) Interest earned during this time period = $750 – 600

Interest earned during this time period = $150.

*****